Help me please!!!
A new machine can make 10,000 aluminum cans three times faster than an older machine. With both machines working, 10,000 cans can be made in 9 hours. How long would it take the new machine, working alone, to make the 10,000 cans?
I don't know how to show my work on this! :(

Answer by stanbon(75887)

A new machine can make 10,000 aluminum cans three times faster than an older machine. With both machines working, 10,000 cans can be made in 9 hours. How long would it take the new machine, working alone, to make the 10,000 cans?------------------------------------
Older machine DATA:
time = 3x hrs/job ; rate = 1/3x jobs/hr
--------
Newer machine DATA:
time = x hrs/job ; rate = 1/x job/hr
-------------
Together DATA:
time = 9 hrs/job ; rate = 1/9 job/hr
=======================================
Equation:
rate + rate = together rate
1/3x + 1/x = 1/9
Multiply thru by 9x to get:
3 + 9 = x
---
x = 12 hrs (time for the newer machine)
=======================
Cheers,
Stan H.